A class of 50 students obtained the following grades in a statistics course:
A 5 B 25 C 15 D 5
If one student is selected at random, what is the probability that
a. the student obtained a B?
b. the student obtained either a B or a C?
c. the student did not obtain a D?
d. the student did not obtain either a B or a C?

Answer by rothauserc(4718)

a) Probability selected student obtained a B = 25/50 = 1/2
:
b) Probability selected student obtained a B or a C = 25/50 + 15/50 = 40/50 = 4/5
:
c) Probability selected student did not obtain a D = 1 - 5/50 = 45/50 = 9/10
:
d) Probability selected student did not obtain a B or a C = 1 - 4/5 = 1/5
: